Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News

  • Despite some non-agreeable weather (winds!), high frequency observing over the weekend! Data for several projects obtained (K, Ka, and Q) that need to be reviewed quickly.
  • OST under operations.
  • Bob, Steve, Steve: Back from International SKA Science and Engineering Meeting.

Commissioning review/reports

Notes for week of 29 Mar 2010 (Black: old but still relevant; Blue: new)

  • Went to a new system for the status of the OSRO observing/priorities. See:
  • Priorities notes (Commissioning, OSRO, RSRO, Demo science)
    • Key drivers:
      • Multi-configuration SBs
      • Push to wider bandwidths
    • Observing priorities
      • Baseline practice:
        • Daytime: Priority for commissioning tests/system verification.
        • Nighttime: Priority for OSRO+ proposals/testing
        • Weekends: Priority for OSRO+ proposals/testing
        • Opportunistic with regard to high frequency observing proposals at high priority, i.e., willing to interrupt, defer maintenance, commissioning test activities if weather allows.

  • Quick Software Systems review:
    • OPT (Lorant):
    • Executor (Barry): Changes and bug fixes (per Barry):
      • 1) Feed a osculating orbit description to CALC.
      • 2) Change port number for vciMapper.
      • 3) Dele one more minor VLA leftover;
      • 4) Make getEopset() static public.
    • Scheduling (Joan, Bryan):
      • Use of API
      • Dynamic scheduling status.
    • CASA (Juergen):
      • Nick in CV this week.
      • Release candidate expected Wednesday.
      • Status of importevla
    • AIPS (Eric): REWAY task
      • New task of interest primarily for EVLA. It attempts to determine the data weights from the rms of the spectral channels, using ICHANSEL to avoid edge effects, RFI, and real line signals. For the weights to be calibrated, the data being fed REWAY must be calibrated or the calibration applied while running REWAY. Note that REWAY can be a substitute for a one source at a time SPLIT. The simple option determines the weight in each time, IF, and polarization product independently. Alternatively, the task can average the data in time and then solve for the weights on an antenna basis, still treating each IF and polarization product independently. These antenna weights may be smoothed in time with a Gaussian before being applied to the data.
  • Hardware notes
    • New switch (between BB and CBE) this week?

OSRO Priority Testing Status

System Verification

  • WIDAR update (Michael)
    • High frequency observing
      • Observing and hand editing
      • Archive
      • Results from the weekend (and what reviewers should look for)
    • Narrow BW (< 500 kHz) status
    • Handling of multi-configuration SBs
    • OSRO2 and zero amplitude visibilities update
    • Missing antennas during observing
  • Array/Receivers
    • High Frequency Readiness
      • See Band checkouts for K, Ka.
      • New reference pointing results. (ECDATA-91): Michael
      • Collimation/TelCal solution issues.
      • Bandpass/broad line checks.
  • Gain transfer: ok
  • OSRO2 BW/channelization Tests (ECDATA-75; Vivek; Reported last Thursday as fine.
  • Anti-aliasing (Michael)
  • Sanity checks
    • Daily checklist
    • Post-maintenance day checklist

Band Checkouts (ECDATA-55)

  • L Band Status: Ready
    • Frequency labeling (ECDATA-9, ECDATA-10): Ok
    • Sky orientation (ECDATA-10): Ok
    • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
    • Tuning: updated. Okay? Still potentially some issues with the new scheme.
      • The "1400 MHz" option: 1328/1456
      • The "1800 MHz" option: 1732/1860
      • The "most likely to be RFI-free" option: 1452/1820
      • The "maximum frequency spread" option: 1348/1860
    • Antennas with no L-band: EA06, EA20, EA26
    • Integrating down (ECDATA-23):
      • Averaging in time is ok; averaging between subbands has problems.
      • L band (really) blank field (ECDATA-40)
    • L band, HI checkout (ECDATA-47); Juergen showed the results of using three tunings to enable adequate bandpass calibration. Will write up recommendation to be passed along.

  • S Band Status: Ready.
    • S band edge tests (similar result to C for tunings near 4 GHz; potentially some satellite downlink RFI; review of aliasing/out-of-band intermodulation is under review (Rick)
    • Frequency labeling (ECDATA-24): Ok
    • Sky orientation (ECDATA-18): Ok
    • Gain stability (ECDATA-24): Ok (no jumps seen)
    • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
    • Integrating down (ECDATA-18): Ok
    • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28

  • C Status: Ready
    • Frequency labeling: Ok. Will get more information from Eric's analysis.
    • Sky orientation (ECDATA-35): Ok
    • Gain stability (ECDATA-35): Ok
    • Tuning: Low frequency tuning issue; stepped up for last AH1004.
    • Integrating down (ECDATA-25):
    • Antennas with no C-band: EA13

  • X Band Status: Ready
    • Frequency labeling : ECDATA-17): Ok
    • Sky orientation (ECDATA-17): Ok
    • Gain stability (ECDATA-26): Ok (no jumps)
    • Tuning: No issues
    • Integrating down (ECDATA-26): Ok
    • Antennas
    • RFI: none seen

  • Q Band Status:

Mode testing

  • Polarimetry Status:
    • C (ECDATA-12): Ready
    • L (ECDATA-42): Ready
    • Eric reviewing AH1004 data for AIPS
    • George will be pushing things into JIRA shortly; noted that he is seeing an East/West difference in the data, don't settle to the same value; good candidate for a more detailed discussion (e.g., ionospheric effects, etc).
    • *Rick noted that the design for L, S, and C cross bar will be much better balanced and should show lower cross polarization.
  • Spectral Line
    • Bandpass stability: ECSV-35
    • Bandpass tests: ECSV-92
      • Update from Jacqueline.
  • Wideband L (AB1345 - ) Sanjay

OSRO Proposal Testing (ECDATA-34)

  • AE176 2nd run ([[https://bugs.aoc.nrao.edu/browse/ECDATA-71][ECDATA-71]): Gustaaf, Jacqueline; released.
  • AK730 triggered GRB observations ([[https://bugs.aoc.nrao.edu/browse/ECDATA-72][ECDATA-72]): Joe, Dale; released.
  • AY 199 HI Run 1 ([[https://bugs.aoc.nrao.edu/browse/ECDATA-73][ECDATA-73]): Emmanuel; released.
  • AY 199 HI Run 2(ECDATA-123): Emmanuel
  • AS1012 (ECDATA-74): Juergen; in progress.
  • AU129 (ECDATA-77): Gustaaf; Several released.
  • AR704 - rerun
  • DEMO SCIENCE (ECDATA-88): James. Blocking filler issue.
  • AL733 (ECDATA-122, ECDATA-123): Miriam
  • AS1008 (ECDATA-120): Veronica
  • AR704 (ECDATA-123): Juergen
  • High frequency
    • AC980 (ECDATA-116, ECDATA-123): Manuel
    • AI139 (ECDATA-118): Maurilio
    • AW769 (ECDATA-124): Chris
    • AC974
    • AS1013
    • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • High Priority [Pokorny]: Auto-controlled CBE writes scrambled visibilities.
      • High Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [Butler]: Scan labeling synchronization.
      • Med Priority [Perley]: Issues integrating down when averaging by channel.
      • Med Priority [Myers]: C band 128 MHz spectrum bends down at the edges.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). EA03 R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Also seen on EA07, EA25 (both RR and LL). EA03 should be fixed.
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

Commissioning plans for the week

System Tests

    • [Michael]: Enable multi-configuration SBs (with no hand-editing)
    • [Joan/Michael/Bryan]: Enable Dynamic Scheduling (diagnostics/debugging)
    • [Rick]: Enable RSRO Mode Observing
    • [Rupen; OSRO Priority 0]: OPT-> Archive path complete
    • OSRO-1 (256 MHz): Still need tests of polarization phase convention

OSRO project tests

RSRO

  • Observations
  • Residents
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Adam Deller (POC: Bryan Butler; area(s): Correlator checkout)
    • James Miller-Jones (POC: Claire Chandler; area(s): Demo Science)
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests)
    • Dave Green (arriving; POC: Sanjay Bhatnagar); area(s): Imaging

Demo Science

  • Summer school data plans: 3C391 continuum mosaic (C band)
  • Other projects discussion (Claire)

This week in JIRA

  • No new topics

Actions

  • [van Moorsel]: Outline OSRO observation checklist (all required steps in delivery of archive data to PI).
  • [McMullin]: Add ECSO Projects to JIRA/tracking.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Description of guideline reduction steps in continuum processing.
  • [Perley/Dhawan]: Develop test to develop/demonstrate readiness for high frequency observing.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Perform a frequency sweep test at S band to provide to engineers for review (Q2).
  • [Benson]: Update when automated handling of multiple spectral windows is available (Q2). Clarified: John will organize a small group this week (likely necessary candidates are Eric, Michael, Rick) to resolve the near and long term plans for handling the automated export to UVFITS for heterogeneous spectral windows (e.g., auto split by spectral window, etc). Results to be presented next Monday.

AOB

Planning review (for Q1/Q2 2010):

Note: Beginnings of group pages from the organization page:

Followup deferred. Begin to look at the detailed needs for the tests being scoped out by the leads. Want to finalize the detailed observing/testing plans by the end of this month and have a view of the longer term commitments needed. Establish an overview of the obstacles to deploying to the community; in particular, understand all of the details of the end-to-end data flow/processing to make decisions on what is working versus what is usable by the community. Something like:

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
Example limit of 1 s int ok OPT selected limits not possible no recommendations avail ok can't cook coffee weather table - transient issue no-go

  • Critical Commissioning
    • System Verification (Perley/Rupen): Need plan to conclude complex gain, bandpass stability tests at L-Q, reference pointing, gain curves, tipping scans.
    • Observing Band Checkout (principally calibration strategies, flagging)
    • Observing Mode Support (broad modes based on science goals)
    • Observing support (e2e capability)
      • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
      • Flux Calibrator Models (Amy); ECSV-43
  • Advanced Science Support

    • Need:
      • [Butler] Planning: Planetary observing (OPT setup; correlator needs); ECSV-7, ECSV-15
      • [Chandler] Planning: Demo Science
      • [Frail] Planning: ToO scheduling heuristics
      • [Sjouwerman] Planning: OPT tests for primary frequencies; ECSV-6
      • [van Moorsel] Planning: Data distribution plan
      • [Sjouwerman] Planning: Archive functionality tests; ECSV-11
      • [Mioduszewski] Planning: Calibrator survey; ECSV-27
      • [Brisken] Planning: Phased array (VLBI, pulsars); ECSV-22, ECSV-23
      • [Golap] Planning CASA multi-thread/cluster testing.
      • [Walker] Planning VLBI setup

-- JosephMcMullin - 29 Mar 2010
Topic revision: r4 - 2010-03-30, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback